Assume that the market for rental cars for business purposes is perfectly competitive, with the demand for this capital input given by K=1500-25v, and the supply is given by K= 74v-500 where K represents the number of cars rented by firms and v is the rental rate per day. What will be the equilibrium levels for v and K in this market?
1
Expert's answer
2020-06-05T10:21:20-0400

v is the rental rate per day 

K is the number of car rented by firms

Demand

K=1500-25v

Supply

K=74v-500

at equilibrium,the quantity demand=the quantity supplied 

inferring that 1500-25v=74v-500

Taking like terms on the same side,we have

1500+500=74v+25v

2000=99v

v=20.20

Demand K=1500-25v

substituting for v

K=1500-25(20.20)

=1500-505.05

K=994